OpenStudy (anonymous):

Hey guys! I would just like to ask if my partial differentiating is correct. f(x,y)=xy (sin (4yx^2)) fyx = sin 4yx^2 + (8yx^2)(cos 4yx^2) + (12yx^2)(cos 4yx^2) - 32(x^4)(y^2)(sin 4yx^2) Thanks!

OpenStudy (ash2326):

You have differentiated with respect to x or y?

OpenStudy (anonymous):

Well, fyx. So I first diff with respect to y then with respect to x. :)

OpenStudy (turingtest):

it's the mixed derivative? let me check...

OpenStudy (ash2326):

Ok :)

OpenStudy (turingtest):

\[f_x=y\sin(4yx^2)+8x^2y^2\sin(4x^2y)\cos(4x^2y)\]I'm typing it out so we can all see if I make a mistake...

OpenStudy (anonymous):

Fy first. :) Then diff Fy with respect to x.

OpenStudy (turingtest):

the order shouldn't matter I don't think\[f_{xy}=\sin(4yx^2)+4yx^2\cos(4yx^2)+\frac{\partial}{\partial y}[8x^2y^2\sin(4x^2y)\cos(4x^2y)]\]

OpenStudy (anonymous):

Actually in some cases the order does not matter. But in others, it does. Anyway, thank you turing test! :)

OpenStudy (turingtest):

oh but I have a 16...

OpenStudy (turingtest):

I'll do it the other way on paper to see if that makes a diff

OpenStudy (turingtest):

ok I redid it the other way and got the same answer as you sorry for the confusion

OpenStudy (anonymous):

It's okay Turing Test. I'm sorry for taking some of your time. :))) But, thanks again. I appreciate your effort.

OpenStudy (turingtest):

no it's my fault for messing up the first time you're welcome